# Designing reaction-cross-diffusion systems with Turing and wave instabilities

### v1

@@ 3.2 Wave instabilities
We next prove an analogous result as in the previous lemma for the case of wave instabilities. Due to these requiring a 3-component subsystem, the constructions are a bit more involved.
@@
Proof.
As $\mathbf{\hat{D}}$ is a non-scalar diagonalizable matrix with positive real eigenvalues and the elements in $\mathbb{D}$ are sorted, then there must exist at least one integer $21\leq\ell\leq n-$ 2 such that $\sigma_{\ell}<\sigma_{\ell+2}$.
@@
Now, when we consider the full system, we note that (by the construction of $\mathbf{C}$) $n-3$ eigenvalues of $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}-\mu\mathbb{D}$ have the form $-1-\sigma_{i}$, for a corresponding integer $1\leq i\leq n$ such that $i\neq\ell,\ell+1,\ell+2$. Moreover, the other three eigenvalues are solutions to the following equation:
@@
$\quad+\left(2b-\mu\sigma_{ell+1}\right)\left(b\mu\sigma_{\ell+2}+cy\right).$
@@
which is positive as we have that $0<\sigma_{\ell}\leq\sigma_{\ell+1}\leq\sigma_{\ell+2}$ and $\sigma_{\ell}<\sigma_{\ell+2}$. This implies that the root with a positive real part we had found for (10) is real for a sufficiently large value of $b$, concluding that there exists a positive real value of $\mu$ such that $\lambda(\mu)=\beta+i\omega$, where $\beta$ can be taken to be positive, thus ensuring the existence of a wave instability. ∎
@@
Proof of Theorem 2.
The proof is analogous to the proof of Theorem 1, though simpler as we can directly use Lemmas 4 and 5 to design unstable submatrices of $\mathbf{J}$ fulfilling the criteria for each kind of instability. ∎
We remark that, as long as one ensures that $\mathbf{J}$ is a stable matrix, there is otherwise freedom to design all other nonzero elements of the matrix away from those that ensure these specific unstable submatrices. With more effort, one can modify other elements of $\mathbf{J}$, but will need to use more general constructions than those proposed here.
@@ 4 Examples
We now employ the constructions used in the proofs of the main theorems, as well as some additional results collected in Appendix A to illustrate the design of Turing and wave instabilities in a handful of prototypical reaction-cross-diffusion systems. The examples here demonstrate the key ingredients needed to use the results, and how the freedom in the constructions allows one to satisfy additional constraints and still achieve the correct instabilities.
@@
[…]
which has eigenvalues $-1$ and $\frac{1}{24}\left(-31\pm\sqrt{335}i\right)$. Table 1 shows the eigenvalues of the principal submatrices of $\mathbf{J}$.
@@
$\mathbf{S_{1}}$ $-1/3$
$\mathbf{S_{2}}$ $-9/4$
$\mathbf{S_{3}}$ $-1$
@@
$\mathbf{S_{1,3}}$ $-1,-1/3$
$\mathbf{S_{2,3}}$ $-1,-9/4$
Table 1: Eigenvalues of the principal submatrices of $\mathbf{J}$ for the 3-component Schnakenberg system (11) when $(a,b)=(1,0.5)$.
@@
Hence, for sufficiently small $\delta>0$, we anticipate that the full system exhibits a wave instability. To ensure that the pattern is visible, i.e., that there is a nearby stable patterned state, we compute the criticality of both bifurcations, the traveling wave and the standing wave, using the approach outlined in Villar-sepulveda-wave . The bifurcation for both, travelling and standing waves turn out to be supercritical in the set of values $(b,\delta)\in[0,1]\times[0,1]$ on an infinite domain, indicating that they should be feasible to obtain via simulations. However, only the traveling wave turns out to be a stable pattern with periodic boundary conditions (see knobloch ), whilst the standing wave turns out to be stable in the presence of homogeneous Neumann boundary conditions. In particular, for $\delta=0.8$ we set the length of the domain as $L=2.38\pi$ and integrated the system with periodic boundary conditions. The solution we obtained is shown in Figure 3, where we see that a travelling wave emerges from a smaller amplitude standing wave transient. We also integrated the system with homogeneous Neumann boundary conditions to observe the solution shown in Figure 4.
Figure 3: Solution $u$ of the 3-component Schnakenberg system (11) with periodic boundary conditions, and with the diffusion matrix given by (15) and $(a,b,\delta)=(1,0.5,0.8)$.
Figure 4: Solution $u$ of the 3-component Schnakenberg system (11) with homogeneous Neumann boundary conditions, and with a diffusion matrix given by (15) and $(a,b,\delta)=(1,0.5,0.8)$.
@@ 4.2 Hyperbolic reaction-diffusion system
We next provide an example based on a hyperbolic system. In Andrew-hyperbolic , the authors study a general class of second order in time cross-diffusion systems. Such systems have particular motivations coming from understanding the microscopic inertia responsible for preventing the infinite speed of propagation that occurs in usual diffusion equations. They also arise as truncations or linearizations of some nonlocal-in-time systems, e.g. those due to fixed or distributed time delays arising in gene expression sargood2022fixed . As an example of applying our constructions to these systems, we consider the following system from Andrew-hyperbolic :
@@
While the analysis in Andrew-hyperbolic can be used to determine if a given set of parameters admits Turing or wave instabilities, the approach we use below allows us to constructively build a particular set of diffusion parameters (for a slightly modified system). Importantly, the method here generalizes to systems of $n>2$ hyperbolic equations, though it is easier to see in this simplified setting.
@@
[…]
@@
respectively. This implies that we can make any of these matrices unstable and that will let us generate diffusion-driven-instabilities through $\sigma_{1}$ or $\sigma_{2}$, respectively. For concreteness, we set $\sigma_{2}=1$ and leave $\sigma_{1}$ as a free non-negative parameter.
@@
which has positive real eigenvalues. We then take $\sigma_{1}=\delta$ and vary this parameter to generate the dispersion relation shown in Figure 5, where $\delta$ goes from 0 to 2 in uniform steps of size 0.4. This graph is particularly interesting as it lets us see that, depending on the value of $\delta$, we find an interaction between Turing and wave instabilities (the eigenvalue with the largest real part is complex with a non-zero imaginary part for low values of $k$ but it becomes real for larger wavenumbers). Note that, for each $\delta$ and $k$, we are plotting the two eigenvalues with the largest real parts, which each coincide with the complex-conjugate pairs for small $k$ but bifurcate for larger $k$ into two strictly real branches. As before, the orange curves correspond to the inadmissible value of $\delta=0$, at which each curve tends to a different asymptotic eigenvalue shown in green lines. Again, we obtain the correct instability for those curves between the orange and red curves.
In particular, if we set homogeneous Neumann boundary conditions, take $\delta=1.3$ with a domain length $L=6$, and integrate the system, we obtain the pattern shown in Figure 6. On the other hand, if we set periodic boundary conditions, with the same value of $\delta$, but a domain length $L=30$ and integrate the system, we obtain the pattern shown in Figure 7.
@@
Figure 6: Solution $u$ of the 2-component hyperbolic system (16) with homogeneous Neumann boundary conditions, a diffusion matrix given by (17), $(a,b,\tau)=(0.257,0.98,0.1)$, and the diffusion parameters provided in (18), together with $\left(\sigma_{1},\sigma_{2}\right)=(1.3,1)$.
Figure 7: Solution $u$ of the 2-component hyperbolic system (16) with periodic boundary conditions, a diffusion matrix given by (17), $(a,b,\tau)=(0.257,0.98,0.1)$, and the diffusion parameters provided in (18), together with $\left(\sigma_{1},\sigma_{2}\right)=(1.3,1)$.
@@
The variables $H$ and $I$ represent the populations of healthy and infected humans, respectively, whilst $P$ is the population of infected mosquitoes. Moreover, $b_{H}$ represents the birth rate of healthy humans, $d_{H}$ the death rate of humans, $c$ the infection rate between humans, $r$ the human recovery rate, $d_{M}$ the death rate of infected mosquitoes, and $b$ is the infection rate between mosquitoes (see alonso for discussion of substantially more complicated compartment-based models of malaria transmission dynamics, from which the present model is a simplified albeit spatially-extended variant). Note that healthy mosquitoes do not appear explicitly in this model, as $Q$ represents the total (fixed) mosquito population, so $Q-P$ is the population of healthy mosquitoes.
@@
[…]
@@
we find that $M_{3}=3.32887$ when $\delta=0$, so this can act as a principal unstable submatrix. Moreover, when we let $\delta$ range from 0 to 0.1 in discrete steps, we obtain the dispersion relations shown in Figure 8.
@@
$t=200$                   $t=500$                   $t=1000$
Figure 9: Three time points of the $H$-component of the solution of the 3-component malaria model (19) solved in a square domain of side length $L=6$ with homogeneous Neumann boundary conditions. The diffusion matrix is given by (22) with the kinetic parameters and diffusion rates as in (23) and (24), respectively. Solutions in $P$ and $I$ are out-of-phase with these patterns, but otherwise qualitatively similar.
We show examples of two-dimensional plots obtained in Fig. 9, where the small perturbation of the homogeneous equilibrium relaxes to a stationary pattern over time. We remark that the choice of nonlinear diffusion here was motivated in part because linear cross-diffusion terms will not ensure that the variables remain positive, breaking both the boundedness of the solutions due to the quadratic nonlinearities, and the physical interpretations of the state variables as populations. Nevertheless, the theory developed here is immediately applicable to suitable parameterized forms of nonlinear diffusion as shown in this example.
@@ 4.4 Keller-Segel model
This entails that, to make this matrix stable, we need to have $b>0$, $e+j<0$, and $ej-fh>0$. Now, we fix $a=3$. With this, we have two cases:
@@
In this case, following Lemma 9, as there is only one entry in $\mathbb{D}$ with the lowest value located in the lower right corner, we have that the eigenvalues of $\mathbf{J}-\mu\mathbf{\hat{D}}$ will be lead by the bottom right entry of $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$ when said entry scales large. For example, if we consider $d=0.9$, and consider the parameters
@@
we have that the bottom right element of $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$ equals $1100>0$. This gives rise to the dispersion relation shown in Figure 10. This shows that we have generated a diffusion-driven instability for the parameter values chosen. Furthermore, when $k=5$, we have that $\lambda_{\max}=20.4439\in\mathbb{R}$, which implies that we are seeing a Turing instability.
Furthermore, to check our prediction, we integrated the system numerically and obtained the solution shown in Figure 11. The solution initially forms a series of spikes that merge, eventually tending towards peaks commonly seen in Keller-Segel type models more generally horstmann20031970 ; painter2011spatio .
Figure 10: Dispersion relation for the Keller-Segel model (25) with (26), for $a=3$ and the parameter values (28), with $d=0.9$.
Figure 11: Solution $u$ of the Keller-Segel model (25) with (26), when the parameters are fixed as in (28) with $a=3$ and $d=0.9$.
@@
which is unstable with two complex conjugate eigenvalues. This gives rise to the dispersion relation shown in Figure 12, which shows that we have generated a diffusion-driven instability for the parameter values chosen. Furthermore, when $k=4$, we have that $\lambda_{\max}=26.1765\pm 75.7567i\in\mathbb{C}$, which implies that we are seeing a wave instability.
Once again, to check the outcome of the design process we have carried out, we integrated the system with homogeneous Neumann boundary conditions for these parameter values. In Figure 13, we show a solution leading to spatiotemporal oscillations, which is an outcome one expects from a wave instability Villar-sepulveda-wave .
Figure 12: Dispersion relation for the Keller-Segel model (25) with (26), for $a=3$ and the parameter values (29), with $d=1.1$.
Figure 13: Solution $u$ of the Keller-Segel model (25) with (26) using homogeneous Neumann boundary conditions, when the parameters are fixed as in (29) with $a=3$ and $d=1.1$.

### v2

@@ 3.2 Wave instabilities
We next prove an analogous result to the previous lemma for the case of wave instabilities. Due to these requiring a 3-component subsystem, the constructions are a bit more involved.
@@
Proof: As $\mathbf{\hat{D}}$ is a non-scalar diagonalizable matrix with positive real eigenvalues and the elements in $\mathbb{D}$ are sorted, then there must exist at least one integer $1\leq\ell\leq n-2$ such that $\sigma_{\ell}<\sigma_{\ell+2}$.
@@
Now, when we consider the full system, we note that (by the construction of $\mathbf{C}$) n3n-3 eigenvalues of $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}-\mu\mathbb{D}$ have the form $1-\sigma_{i}$, for a corresponding integer $1\leq i\leq n$ such that $i\neq\ell,\ell+1,\ell+2$. Moreover, the other three eigenvalues are solutions to the following equation:
@@
$\quad+\left(2b-\mu\sigma_{\ell+1}\right)\left(b\mu\sigma_{\ell+2}+cy\right).$
@@
which is positive as we have that $0<\sigma_{\ell}\leq\sigma_{\ell+1}\leq\sigma_{\ell+2}$ and $\sigma_{\ell}<\sigma_{\ell+2}$. This implies that the root with a positive real part we had found for Eq. 10 is real for a sufficiently large value of $b$, concluding that there exists a positive real value of $\mu$ such that $\lambda(\mu)=\beta+i\omega$, where $\beta$ can be taken to be positive, thus ensuring the existence of a wave instability.
@@
Proof: [Proof of Theorem 2] The proof is analogous to the proof of Theorem 1, though simpler as we can directly use Lemmas 4 and 5 to design unstable submatrices of $\mathbf{J}$ fulfilling the criteria for each kind of instability. We remark that, as long as one ensures that $\mathbf{J}$ is a stable matrix, there is otherwise freedom to design all other nonzero elements of the matrix away from those that ensure these specific unstable submatrices. With more effort, one can modify other elements of $\mathbf{J}$, but will need to use more general constructions than those proposed here.
@@ 4 Examples
We now employ the constructions used in the proofs of the main theorems, as well as some additional results collected in Appendix A to illustrate the design of Turing and wave instabilities in a handful of prototypical reaction-cross-diffusion systems. The examples here demonstrate the key ingredients needed to use the results and how the freedom in the constructions allows one to satisfy additional constraints and still achieve the correct instabilities.
@@
$\mathbf{J}=\left(\begin{array}[]{ccc}-\frac{1}{3}&\frac{9}{4}&0\\[4.30554pt] -\frac{2}{3}&-\frac{9}{4}&0\\[4.30554pt] 1&0&-1\end{array}\right),$
which has eigenvalues 1-1 and $\frac{1}{24}\left(-31\pm\sqrt{335}i\right)$. Table 1 shows the eigenvalues of the principal submatrices of $\mathbf{J}$.
@@
$\mathbf{S_{1}}$ 1/3-1/3
$\mathbf{S_{2}}$ 9/4-9/4
$\mathbf{S_{3}}$ 1-1
@@
$\mathbf{S_{1,3}}$ 1,1/3-1,-1/3
$\mathbf{S_{2,3}}$ 1,9/4-1,-9/4
Table 1: Eigenvalues of the principal submatrices of $\mathbf{J}$ for the 3-component Schnakenberg system, Eq. 11, when $(a,b)=(1,0.5)$.
@@
Hence, for sufficiently small $\delta>0$, we anticipate that the full system exhibits a wave instability. To ensure that the pattern is visible, i.e., that there is a nearby stable patterned state, we compute the criticality of both bifurcations, the traveling wave and the standing wave, using the approach outlined in Villar-sepulveda-wave . The bifurcation for both, travelling and standing waves turns out to be supercritical in the set of values $(b,\delta)\in[0,1]\times[0,1]$ on an infinite domain, indicating that they should be feasible to obtain via simulations. However, only the traveling wave turns out to be a stable pattern with periodic boundary conditions (see knobloch ), whilst the standing wave turns out to be stable in the presence of homogeneous Neumann boundary conditions. In particular, for $\delta=0.8$ we set the length of the domain as $L=2.38\pi$ and integrated the system with periodic boundary conditions. The solution we obtained is shown in Fig. 3, where we see that a travelling wave emerges from a smaller amplitude standing wave transient. We also integrated the system with homogeneous Neumann boundary conditions to observe the solution shown in Fig. 4.
Figure 3: Solution $u$ of the 3-component Schnakenberg system, Eq. 11, with periodic boundary conditions, and with the diffusion matrix given by Eq. 15 and $(a,b,\delta)=(1,0.5,0.8)$.
Figure 4: Solution $u$ of the 3-component Schnakenberg system, Eq. 11, with homogeneous Neumann boundary conditions, and with a diffusion matrix given by Eq. 15 and $(a,b,\delta)=(1,0.5,0.8)$.
@@ 4.2 Hyperbolic reaction-diffusion system
We next provide an example based on a hyperbolic system. In ritchie2022turing , the authors study a general class of second order in time cross-diffusion systems. Such systems have particular motivations coming from understanding the microscopic inertia responsible for preventing the infinite speed of propagation that occurs in usual diffusion equations. See mendez2010reaction for an overview of hyperbolic and other formulations of reaction-transport processes aiming to overcome this difficulty, and some of their connections with thermodynamic principles. They also arise as truncations or linearizations of some nonlocal-in-time systems, e.g. those due to fixed or distributed time delays arising in gene expression sargood2022fixed . As an example of applying our constructions to these systems, we consider the following system from ritchie2022turing :
@@
While the analysis in ritchie2022turing can be used to determine if a given set of parameters admits Turing or wave instabilities, the approach we use below allows us to constructively build a particular set of diffusion parameters (for a slightly modified system). Importantly, the method here generalizes to systems of $n>2$ hyperbolic equations, though it is easier to see in this simplified setting.
@@
$u^{*}$ $=\frac{\sqrt[3]{2}M^{2}+6a^{3}-2a^{2}}{3\cdot 2^{2/3}a^{2}M},$
$v^{*}$ $=\frac{-2^{2/3}M^{2}-6\sqrt[3]{2}a^{3}+2a^{2}\left(3bM+\sqrt[3]{2}\right)}{18a^{3}M},$
@@
respectively. This implies that we can make any of these matrices unstable, and that will let us generate diffusion-driven instabilities through $\sigma_{1}$ or $\sigma_{2}$, respectively. For concreteness, we set $\sigma_{2}=1$ and leave $\sigma_{1}$ as a free non-negative parameter.
@@
which has positive real eigenvalues. We then take $\sigma_{1}=\delta$ and vary this parameter to generate the dispersion relation shown in Fig. 5, where $\delta$ goes from 0 to 2 in uniform steps of size 0.4. This graph is particularly interesting as it lets us see that, depending on the value of $\delta$, we find an interaction between Turing and wave instabilities (the eigenvalue with the largest real part is complex with a non-zero imaginary part for low values of $k$ but it becomes real for larger wavenumbers). Note that, for each $\delta$ and $k$, we are plotting the two eigenvalues with the largest real parts, which each coincide with the complex-conjugate pairs for small $k$ but bifurcate for larger $k$ into two strictly real branches. As before, the orange curves correspond to the inadmissible value of $\delta=0$, at which each curve tends to a different asymptotic eigenvalue shown in green lines. Again, we obtain the correct instability for those curves between the orange and red curves.
In particular, if we set homogeneous Neumann boundary conditions, take $\delta=1.3$ with a domain length $L=6$, and integrate the system, we obtain the pattern shown in Fig. 6. On the other hand, if we set periodic boundary conditions, with the same value of $\delta$, but a domain length $L=30$ and integrate the system, we obtain the pattern shown in Fig. 7.
@@
Figure 6: Solution $u$ of the 2-component hyperbolic system, Eq. 16, with homogeneous Neumann boundary conditions, a diffusion matrix given by Eq. 17, $(a,b,\tau)=(0.257,0.98,0.1)$, and the diffusion parameters provided in Eq. 18, together with $\left(\sigma_{1},\sigma_{2}\right)=(1.3,1)$.
Figure 7: Solution $u$ of the 2-component hyperbolic system, Eq. 16, with periodic boundary conditions, a diffusion matrix given by Eq. 17, $(a,b,\tau)=(0.257,0.98,0.1)$, and the diffusion parameters provided in Eq. 18, together with $\left(\sigma_{1},\sigma_{2}\right)=(1.3,1)$.
@@
The variables $H$ and $I$ represent the populations of healthy and infected humans, respectively, whilst $P$ is the population of infected mosquitoes. Moreover, $b_{H}$ represents the birth rate of healthy humans, $d_{H}$ the death rate of humans, $c$ the infection rate between humans, $r$ the human recovery rate, $d_{M}$ the death rate of infected mosquitoes, and $b$ is the infection rate between mosquitoes (see alonso for discussion of substantially more complicated compartment-based models of malaria transmission dynamics, from which the present model is a simplified albeit spatially-extended variant). Note that healthy mosquitoes do not appear explicitly in this model, as $Q$ represents the total (fixed) mosquito population, so QPQ-P is the population of healthy mosquitoes.
@@
This system has only two equilibria, $\mathbf{0}=(0,0,0)$ and $\mathbf{P}=\left(H^{*},I^{*},P^{*}\right)$, where
@@
we find that $M_{3}=3.32887$ when $\delta=0$, so this can act as a principal unstable submatrix. Moreover, when we let $\delta$ range from 0 to 0.1 in discrete steps, we obtain the dispersion relations shown in Fig. 8.
@@
$t=20$         $t=50$         $t=100$
Figure 9: Three time points of the $H$-component of the solution of the 3-component malaria model, Eq. 19, solved in a square domain of side length $L=6$ with homogeneous Neumann boundary conditions. The diffusion matrix is given by Eq. 22 with the kinetic parameters and diffusion rates as in Eqs. 23 and 24, respectively. Solutions in $P$ and $I$ are out-of-phase with these patterns, but otherwise qualitatively similar.
We show examples of two-dimensional plots obtained in Fig. 9, where the small perturbation of the homogeneous equilibrium relaxes to a stationary pattern over time. We remark that the choice of nonlinear diffusion here was motivated in part because linear cross-diffusion terms will not ensure that the variables remain positive, breaking both the boundedness of the solutions due to the quadratic nonlinearities and the physical interpretations of the state variables as populations. Nevertheless, the theory developed here is immediately applicable to suitable parameterized forms of nonlinear diffusion as shown in this example.
@@ 4.4 Keller-Segel model
This entails that, to make this matrix stable, we need to have $b>0$, $e+j<0$, and ejfh>0ej-fh>0. Now, we fix $a=3$. With this, we have two cases:
@@
In this case, following Lemma 9, as there is only one entry in $\mathbb{D}$ with the lowest value located in the lower right corner, we have that the eigenvalues of $\mathbf{J}-\mu\mathbf{\hat{D}}$ will be led by the bottom right entry of $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$ when said entry is large. For example, if we consider $d=0.9$, and consider the parameters
@@
we have that the bottom right element of $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$ equals $1100>0$. This gives rise to the dispersion relation shown in Fig. 10. This shows that we have generated a diffusion-driven instability for the parameter values chosen. Furthermore, when $k=5$, we have that $\lambda_{\max}=20.4439\in\mathbb{R}$, which implies that we are seeing a Turing instability.
Furthermore, to check our prediction, we integrated the system numerically and obtained the solution shown in Fig. 11. The solution initially forms a series of spikes that merge, eventually tending towards peaks commonly seen in Keller-Segel type models more generally horstmann20031970 ; painter2011spatio .
Figure 10: Dispersion relation for the Keller-Segel model, Eq. 25, with (26), for $a=3$ and the parameter values as in Eq. 28, with $d=0.9$.
Figure 11: Solution $u$ of the Keller-Segel model, Eq. 25, with (26), when the parameters are fixed as in Eq. 28 with $a=3$ and $d=0.9$.
@@
which is unstable with two complex conjugate eigenvalues. This gives rise to the dispersion relation shown in Fig. 12, which shows that we have generated a diffusion-driven instability for the parameter values chosen. Furthermore, when $k=4$, we have that $\lambda_{\max}=26.1765\pm 75.7567i\in\mathbb{C}$, which implies that we are seeing a wave instability.
Once again, to check the outcome of the design process we have carried out, we integrated the system with homogeneous Neumann boundary conditions for these parameter values. In Fig. 13, we show a solution leading to spatiotemporal oscillations, which is an outcome one expects from a wave instability Villar-sepulveda-wave .
Figure 12: Dispersion relation for the Keller-Segel model, Eq. 25, with (26), for $a=3$ and the parameter values as in Eq. 29, with $d=1.1$.
Figure 13: Solution $u$ of the Keller-Segel model, Eq. 25, with (26) using homogeneous Neumann boundary conditions, when the parameters are fixed as in Eq. 29 with $a=3$ and $d=1.1$.